# Curvature-Constrained Vector Field for Motion Planning of Nonholonomic Robots

### v1

@@ II-C Proposed Algorithm Framework
To address the motion planning problem formulated above, the algorithm we will propose in this paper has the feedback structure as illustrated in Figure 1. Within this framework, the VF presented in Section III aims to provide the nonholonomic robot with a reference signal including the reference orientation and curvature-bounded reference trajectories that converge to the target positive limit set, and the control laws in Section IV intend to ensure that the kinematically constrained robot tracks the VF with bounded control inputs. By this means, the robot’s orientation aligns with the reference orientation and the robot’s trajectory converges to the target positive limit set by moving along the reference trajectory.
@@ III Curvature-Constrained Vector Field
In this section, we will construct the curvature-constrained vector field (CVF), where a stable limit cycle is established as the target positive limit set. To guide the robot with curvature-bounded reference trajectory, we further provide the condition to ensure that the integral curves of CVF have a prescribed upper bound on the curvature while converging to the limit set.
Before moving forward, we present several concepts about the VFs.
@@
[…]
The curvature of integral curves is given in the following lemma.
@@
Consider a VF $\mathbf{F}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ given in the polar coordinates $(r,\varphi)$. The curvature of the integral curves at the nonsingular point $\bm{p}$ is given by
@@
Intuitively, the VF is naturally advantageous in nonholonomic robot motion planning, since the VF assigns a vector to each position in the workspace, designating the desired velocity direction of the robot. When the integral curves of VF converge to the target positive limit set, the robot arrives at the target by aligning its orientation with the VF and moving along the integral curves. In this sense, the integral curve can be regarded as the reference trajectory for the robot. However, when curvature of the integral curves exceeds the upper bound of the robot’s trajectory curvature, imposed by the curvature constraint, the robot would not be able to follow such reference trajectories. Therefore, the integral curves should not only converge to the target positive limit set but also have prescribed bound on the curvature.
To construct such VF, we utilize three types of flow fields, namely, source, sink and vortex, as illustrated in Figure 2. For simplicity, we assume that the origin is the singular point for all these VFs. Since the VF necessarily specifies the orientation of desired velocity while the magnitude of velocity is determined by control laws, we present the normalized vortex flow field as follows:
$\mathbf{F}_{\text{v}}=\left\{\begin{matrix}&\left[F_{\text{v},r},F_{\text{v},\varphi}\right]^{\intercal}=\left[0,1\right]^{\intercal},&\bm{p}\neq\bm{0},\\ &\bm{0},&\bm{p}=\bm{0},\end{matrix}\right.$ (6)
where $F_{\text{v},\varphi}=1$ implies the vortex is counterclockwise. Aside from vortex, source and sink are two kinds of flow fields exhibiting opposite behaviors. The integral curves of source radially diverge from the singular point, while the integral curves of sink converge to the singular point. The normalized flow fields of source and sink are denoted as $\mathbf{F}_{\text{out}}$ and $\mathbf{F}_{\text{in}}$, respectively, where
$\mathbf{F}_{\text{out}}=\left\{\begin{matrix}&\left[F_{\text{out},r},F_{\text{out},\varphi}\right]^{\intercal}=\left[1,0\right]^{\intercal},&\bm{p}\neq\bm{0},\\ &\bm{0},&\bm{p}=\bm{0},\end{matrix}\right.$ (7)
$\mathbf{F}_{\text{in}}=\left\{\begin{matrix}&\left[F_{\text{in},r},F_{\text{in},\varphi}\right]^{\intercal}=\left[-1,0\right]^{\intercal},&\bm{p}\neq\bm{0},\\ &\bm{0},&\bm{p}=\bm{0}.\end{matrix}\right.$ (8)
(a) Vortex
(b) Source
(c) Sink
Figure 2: Three types of basic flow fields.
In order to present the positive limit set for the integral curve with dynamics $\dot{\bm{\zeta}}=\mathbf{F}(\bm{\zeta})$, it is intuitive to utilize the flow field of vortex to construct a limit cycle since its integral curves are concentric circular arcs about the singular point. To blend the vortex with sink and source, we divide $\mathbb{R}^{2}$ into four regions with the origin $\bm{p}=\bm{0}$ as the singular point:
@@
$\mathcal{A}_{4}=\{\bm{p}\mid r\geq r_{3}\}.$
In the partition above, the workspace is divided into four regions by three circles with radii $0<r_{1}<r_{2}<r_{3}$, which can be conveniently tuned. Besides, we let $\mathcal{O}_{2}=\left\{r=r_{2}\right\}$ denote the boundary between $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$.
In the following, we aim to establish the circle $\mathcal{O}_{2}$ as a stable limit cycle. Inspired by the phase portrait of stable limit cycle in nonlinear dynamics, the evolution of trajectories inside the limit cycle can be decomposed into two orthogonal components: one is outward component radially approaching the limit cycle and the other is circular component parallel to the limit cycle. The effect of these two components can be approximated by the source and the vortex, respectively. Similarly, we can approximate the evolution of trajectories outside the limit cycle by blending the sink and the vortex. To blend the aforementioned flow fields, we employ the blending function:
which smoothly decreases from 1 to 0 on the interval $s\in[s_{l},s_{u}]$ with zero derivatives on both sides. For convenience, we denote $\tilde{\lambda}(s;s_{l},s_{u})=1-\lambda(s;s_{l},s_{u})$.
Based on the blending function, we present the blended flow fields of the vortex with the source and sink for regions $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$ as follows:
@@
Then, we embed the flow fields of source and sink into the regions $\mathcal{A}_{1}$ and $\mathcal{A}_{4}$, respectively, leading to the blended VF $\mathbf{F}$ given by
$\mathbf{F}(\bm{p})=\left\{\begin{matrix}&\mathbf{F}_{\text{out}},&\bm{p}\in\mathcal{A}_{1},\\ &\mathbf{F}_{\text{out}}^{\text{v}},&\bm{p}\in\mathcal{A}_{2},\\ &\mathbf{F}_{\text{v}}^{\text{in}},&\bm{p}\in\mathcal{A}_{3},\\ &\mathbf{F}_{\text{in}},&\bm{p}\in\mathcal{A}_{4}.\end{matrix}\right.$ (13)
In the above construction, we establish the stable limit cycle with three parameters $r_{1}$, $r_{2}$ and $r_{3}$ that can be conveniently adjusted. Based on these parameters, we derive a sufficient condition ensuring that the integral curves of the blended VF $\mathbf{F}$ have bounded curvature while converging to the limit cycle.
@@
The VF $\mathbf{F}$ defined in (13) satisfies the following properties:
@@
    The integral curves of $\mathbf{F}$ have continuous curvature bounded by $\bar{\kappa}$ if there hold that
$r_{i}\geq r_{i+1}-r_{i},\quad i=1,2,$ (14a)
$r_{i+1}-r_{i}\geq 3\rho,\quad i=1,2.$ (14b)
@@
After the proposition of $\mathbf{F}$, the existence of the stable limit cycle $\mathcal{O}_{2}$ for $\mathbf{F}$ inspires us that we can find a point $\bm{p}_{c}\in\mathcal{O}_{2}$ such that the tangent vector $\mathbf{F}(\bm{p}_{c})$ aligns with the target orientation, i.e., $\mathbf{F}(\bm{p}_{c})=[\cos\theta_{d},\sin\theta_{d}]^{\intercal}$. Therefore, we can regard the stable limit cycle as the target positive limit set and present the CVF as below:
@@
The CVF given by (15) has the following properties:
@@
Figure 3: The proposed CVF $\mathbf{T}$, where the stable limit cycle, target configuration and the singular point are illustrated.
The proposed CVF is illustrated by Figure 3. In the above theorem, the property (16a) means that the curvature of integral curves of the CVF is continuous and bounded by $\bar{\kappa}$ for any non-singular position. The property (16b) indicates that the integral curves of the CVF converge to a stable limit cycle containing the target position $\bm{p}_{d}$ almost globally. Moreover, the property (16c) ensures that the orientation of the CVF aligns with the target orientation $\theta_{d}$ at $\bm{p}_{d}$. By following the CVF with desired properties above, the nonholonomic robot converges to the target positive limit set with continuous and bounded trajectory curvature.
@@
Compared to existing VF-based motion planners with curvature constraints, the proposed CVF offers several key advantages. In [34], the VF only converges to the target position and depends on empirical parameter tuning. In contrast, our CVF guarantees that curvature-bounded integral curves simultaneously plan both position and orientation, as specified by the explicit condition in (14). While [38] also addresses position and orientation planning, it places the singularity of the curvature-constrained VF at the target position. Unlike this approach, our CVF establishes a stable limit cycle as the target positive limit set, keeping the target configuration away from the singular point and avoiding singularity during convergence. Although [36] utilizes a stable limit cycle, the limited flexibility in its VF design makes it challenging to derive control inputs that both track the VF and maintain the curvature constraint. In contrast, our CVF offers more flexibility in parameter tuning, facilitating the co-design of the CVF and control laws in the subsequent section.
@@ IV Saturated Control Laws with Dynamic Gain
The CVF presented in Theorem 2 provides curvature-bounded reference trajectories converging to the target positive limit set. As pointed out in Section II-C, we will in this section find feasible control inputs of the nonholonomic robot, including the linear velocity $v_{x}$ and angular velocity $\omega$, to track the CVF under curvature constraint.
In accordance with the motion planning objective, we firstly propose the control law with user-defined bounds as follows:
@@
where $v_{\text{min}}$ is the lower bound, $c_{\bm{p}}$, $c_{\theta}$ are positive scalars, and $k_{v}$ is the gain given by $k_{v}=v_{\text{max}}-v_{\text{min}}$ such that the linear velocity is upper bounded by $v_{\text{max}}$. Moreover, $\theta_{e}$ is the orientation error between the robot and CVF given by
@@
and the reference orientation $\theta_{r}$ specified by the CVF is defined as
@@
It can be observed that the linear velocity $v_{x}$ decreases to the lower bound $v_{\text{min}}$ only if the robot reaches the target configuration $\bm{g}_{d}$, where $\theta_{r}=\theta_{d}$ and hence $\|\bm{p}-\bm{p}_{d}\|=\left|\theta_{e}\right|=0$. For robots such as wheeled UGV, $v_{\text{min}}$ could be chosen as zero so that the robot could asymptotically converge to $\bm{g}_{d}$. As for fixed wing UAVs and other robots with $v_{\text{min}}>0$, they are not able to converge to a certain configuration and thereby the result of convergence to the target positive limit set is the best we can achieve.
When the robot’s orientation $\theta$ aligns with the reference orientation $\theta_{r}$, the robot’s trajectory becomes the time reparameterization of the integral curves of the CVF under the proposed linear velocity, referring to [39, Chapter 1.5]. This means that the robot is able to track the CVF and converge to the target positive limit set by aligning its orientation with $\theta_{r}$ under the angular velocity:
@@
In the above equation, $-k_{\omega}\theta_{e}$ is the feedback term and the feedforward term, i.e., the time derivative of $\theta_{r}$, is given by
@@
where $\Delta\theta$ is the difference between the robot’s orientation $\theta$ and the orientation of the gradient $\nabla\theta_{r}$, the magnitude of gradient is denoted by the amplitude $A(r_{\delta})=\|\nabla\theta_{r}\|$ and $(r_{\delta},\varphi_{\delta})$ is the polar coordinate centering at the singular point $\bm{p}_{\delta}$ defined as
$r_{\delta}$ $=\|\bm{p}-\bm{p}_{\delta}\|,$
[…]
Detailed derivations of (21) are provided in Appendix D.
To ensure that the robot has curvature-bounded trajectory before it tracks the CVF, we refer to (2) and impose the angular velocity with a saturation bound defined as
@@
and hence the saturated angular velocity is given by
@@
where $\omega_{0}$ is given by (20), $\mathrm{sgn}(\cdot)$ is the sign function. According to (43) and (44) in Appendix D, there holds that
@@
The equations above implies that the upper bound of the feedforward term is larger than the saturation bound when $\bm{g}\in\mathcal{S}=\{\bm{g}\mid\|\bm{p}-\bm{p}_{\delta}\|<\rho\}$. Therefore, the angular velocity (22) possibly saturates in $\mathcal{S}$ even when the feedback term is omitted. In other words, $\mathcal{S}$ is the minimal region where saturation occurs with the presence of feedback term.
To stabilize the orientation error $\theta_{e}$, we scheme to establish the condition under which the angular velocity $\omega$ in (22) does not saturate persistently. In this way, $\theta_{e}$ is eventually stabilized by $\omega=\omega_{0}$. From a perspective of geometry, the robot’s trajectory with curvature bounded by $\bar{\kappa}$ cannot be enclosed in an open circular area with radius $\rho=1/\bar{\kappa}$ such as $\mathcal{S}$. Enlightened by this, we ensure that (22) does not saturate persistently by confining the region where saturation occurs to the minimal extent of $\mathcal{S}$, as elaborated in the following theorem.
@@
The proposition that the angular velocity does not saturate persistently can be verified with a counter example. Assume that the angular velocity saturates all the time, which requires the robot to stay in $\mathcal{S}$. Since the target configuration $\bm{g}_{d}\notin\mathcal{S}$, the linear velocity satisfies $v_{x}>v_{min}\geq 0$. In this case, the robot moves on a circle with radius $\rho$ since $\left|\omega\right|=\bar{\omega}$ and $\kappa=\left|\bar{\omega}/v_{x}\right|=\bar{\kappa}=1/\rho$. However, such a circle cannot be contained within $\mathcal{S}$ since $\mathcal{S}$ itself is an open circular region with radius $\rho$. This implies that the robot will leave $\mathcal{S}$ even when its angular velocity persistently saturates, which contradicts with Theorem 3. Above discussions further imply that the robot leaves the region $\mathcal{S}$ and the saturation only occurs in finite time.
Due to the geometry of the CVF, i.e., the magnitude $A(r_{\delta})$ of gradient $\nabla\theta_{r}$ exceeds the maximum curvature around the singular point $\bm{p}_{\delta}$. Therefore, the saturation of the angular velocity within $\mathcal{S}$ is inevitable. To ensure the robot leaves such region and hence the angular velocity only saturates in finite time, we propose the dynamic gain in the above theorem. However, this dynamic gain vanished when the robot circulates on the boundary of $\mathcal{S}$, that is $k_{\omega}=0$, $\forall\bm{g}\in\mathcal{N}_{k_{\omega}}=\left\{\bm{g}\mid r_{\delta}=\rho,\lvert\theta_{e}\rvert=\pi/2\right\}$. In addition, due to the singularity of the CVF and the definition of orientation error, there are two singular sets for the robot, i.e., $\mathcal{N}_{\mathbf{T}}=\left\{\bm{g}\mid\bm{p}=\bm{p}_{\delta}\right\}$ and $\mathcal{N}_{\theta}=\left\{\bm{g}\mid\lvert\theta_{e}\rvert=\pi\right\}$. Above singular sets together form the non-converging set $\mathcal{N}=\mathcal{N}_{\mathbf{T}}\cup\mathcal{N}_{\theta}\cup\mathcal{N}_{k_{\omega}}$ of the closed-loop system.
It is worth nothing that the dynamic gain not only depends on the robot’s states, but also accounts for the change in CVF under robot’s motion. This strategy bridges the design of CVF and control laws with the equations (24a) and (24b) in dynamic gain. More specifically, the dynamic gain confines the saturation within $\mathcal{S}$ by not only controlling the magnitude of feedback term but also imposing the inherent constraint $A(r_{\delta})\leq\bar{\kappa}$, $\forall r_{\delta}\geq\rho$ on the geometric properties of CVF. To address the above intrinsic coupling between the CVF and control laws under the curvature constraint, the co-design of the CVF and control laws is of great necessity. Thanks to the CVF’s flexible design criterion (14), we are able to conveniently adjust the parameters of CVF.
In the following theorem, we establish the condition under which the orientation error is monotonically stabilized by the saturated control laws with dynamic gain.

### v2

@@ II-C Proposed Algorithm Framework
To address the motion planning problem formulated above, the algorithm to be proposed in this paper employs a feedback structure as illustrated in Figure 1. In this framework, the reference trajectory and reference orientation are provided by the vector field presented in Section III and the control laws to track the vector field are derived in Section IV. Due to the co-design strategy, the vector field adjusts its parameters to ensure the boundedness of control inputs and the control laws updates the feedback gain in response to the change of vector field. In this manner, the robot tracks the vector field and converges to the target positive limit set with prescribed curvature bound.
@@ III Curvature-Constrained Vector Field
In this section, we will construct the curvature-constrained vector field (CVF), where a stable limit cycle that contains the target configuration is established as the target positive limit set. To provide the robot with curvature-bounded reference trajectory, we further derive the condition to ensure that the integral curves of CVF have a prescribed upper bound on the curvature while converging to the limit set.
Before moving forward, we present several concepts about the vector field (VF) [lee2012introduction, Chs. 8, 9].
@@
For simplicity, we consider the polar coordinates $(r,\varphi)$ transformed by $r=\sqrt{x^{2}+y^{2}}$ and $\varphi=\arctan{(y/x)}$. Denote the basis of the Cartesian and polar coordinates as $\left\{\bm{e}_{x},\bm{e}_{y}\right\}$ and $\left\{\bm{e}_{r},\bm{e}_{\varphi}\right\}$, respectively. Then the components of the vector field represented in the polar coordinates are defined as $F_{r}=\left<\bm{e}_{r},\mathbf{F}\right$> and $F_{\varphi}=\left<\bm{e}_{\varphi},\mathbf{F}\right$>, where $\left$$\cdot$$\cdot\right$> is the inner product.
Since the VF designates the desired velocity direction of the robot by assigning a vector to each position in the workspace, the integral curves could be regarded as the reference trajectories for the robot. By aligning the orientation with the VF, the nonholonomic robots’ actual trajectories converge to the integral curves. Therefore, to navigate a curvature-constrained nonholonomic robot, the integral curves should not only converge to the target positive limit set but also have prescribed bound on the curvature. In the following lemma, the computation of curvature of integral curves is shown.
@@
Consider a VF $\mathbf{F}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ given in the polar coordinates $(r,\varphi)$. The curvature of its integral curves at the nonsingular point $\bm{p}$ is given by
@@
To construct a desired CVF with the aforementioned properties, we utilize the elementary flows of vortex, source and sink, which can constitute a wide variety of VFs by superposition[White2016Fluid, Ch. 8]. Specifically, the normalized elementary flows are given, respectively, by the following equations:
$\mathbf{F}_{\text{v}}=\left\{\begin{matrix}&\left[F_{\text{v},r},F_{\text{v},\varphi}\right]^{\intercal}=\left[0,1\right]^{\intercal},&\bm{p}\neq\bm{0},\\ &\bm{0},&\bm{p}=\bm{0}.\end{matrix}\right$. (6)
$\mathbf{F}_{\text{out}}=\left\{\begin{matrix}&\left[F_{\text{out},r},F_{\text{out},\varphi}\right]^{\intercal}=\left[1,0\right]^{\intercal},&\bm{p}\neq\bm{0},\\ &\bm{0}.&\bm{p}=\bm{0}.\end{matrix}\right$. (7)
$\mathbf{F}_{\text{in}}=\left\{\begin{matrix}&\left[F_{\text{in},r},F_{\text{in},\varphi}\right]^{\intercal}=\left[-1,0\right]^{\intercal},&\bm{p}\neq\bm{0},\\ &\bm{0}.&\bm{p}=\bm{0}.\end{matrix}\right$. (8)
Figure 2 shows the above elementary flows, where the integral curves of the vortex are concentric circular arcs around the singular point, while the integral curves of source and sink are straight lines diverging from and converging to the singular point, respectively. The geometric forms of these integral curves match the motion pattern of nonholonomic robots, since the trajectory of such systems can generally be described as compositions of circular arcs and straight-line segments.
(a) Vortex $\mathbf{F}_{\text{v}}$
(b) Source $\mathbf{F}_{\text{out}}$
(c) Sink $\mathbf{F}_{\text{in}}$
Figure 2: Three types of elementary flow with singular point at the origin.
Drawing intuition from the phase portrait of a nonlinear system featuring a stable limit cycle, we construct a desirable VF with the following transition of the elementary flows as the distance from the singular point increases: from the sink near the singularity, to the vortex around the limit cycle, and finally to the source further out. To achieve this transition, we partition the workspace with three concentric circles with radii $0<r_{1}<r_{2}<r_{3}$:
@@
$\mathcal{A}_{4}=\{\bm{p}\mid r\geq r_{3}\},$
where $\mathcal{O}_{2}=\left\{r=r_{2}\right\}$ is targeted as the stable limit cycle with the singular point at the origin. Given the above decomposition, we embed the source and sink in the regions $\mathcal{A}_{1}$ and $\mathcal{A}_{4}$, and then blend the vortex with source and sink in the regions $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$, respectively. Consider a blending function $\lambda(s;s_{l},s_{u})$ that monotonically decreases from 1 to 0 on its domain $s_{l},s_{u}]$ and denote $\tilde{\lambda}(s;s_{l},s_{u})=1-\lambda(s;s_{l},s_{u})$. The blended VF can be given by
$\mathbf{F}(\bm{p})=\begin{cases}\mathbf{F}_{\text{out}},&\bm{p}\in\mathcal{A}_{1},\\ \mathbf{F}_{\text{out}}^{\text{v}},&\bm{p}\in\mathcal{A}_{2},\\ \mathbf{F}_{\text{v}}^{\text{in}},&\bm{p}\in\mathcal{A}_{3},\\ \mathbf{F}_{\text{in}},&\bm{p}\in\mathcal{A}_{4},\end{cases}$ (10)
where the transition between elementary flows is specified by
@@
According to Definition 3 and Lemma 1, the above blended VF has continuous curvature, which is a requirement for the continuity of the tracking control inputs, when the VF is of class $C^{1}$ at least. By calculating the partial derivatives of (10), we can find that the blended VF is of the same class $C^{k}$ as the blending function within $\mathcal{A}_{i=2,3}$. Moreover, the blended VF has continuous partial derivatives on the boundaries between $\mathcal{A}_{i=1,2,3,4}$ only when the blending function has zero derivatives at these boundaries. Therefore, the necessary and sufficient condition for the blended VF to have continuous curvature, is that the blending function $\lambda(r;r_{l},r_{u})$ is of class $C^{k\geq 1}$ and $\frac{\partial\lambda}{\partial r}\mid_{r=r_{l}}=\frac{\partial\lambda}{\partial r}\mid_{r=r_{u}}=0$.
In this paper, we employ the blending function given by
which smoothly decreases from 1 to 0 on the interval $s_{l},s_{u}]$ and has zero derivatives on the boundaries.
It is worth noting that the derivative of this blending function is related to the width of domain. As a result, the transition behavior between different elementary flows in the blended VF can be easily adjusted by modifying the region radii $r_{i}$ as shown in Figure 3. Based on appropriate selection of $r_{i}$, we ensure that the integral curve of the blended VF has a prescribed curvature bound $\bar{\kappa}$ while converging to the limit cycle. This is formally stated in the following theorem.
Figure 3: The blended vector field $\mathbf{F}$, where the domains of source, vortex and sink are shaded in red, blue and green. Within the transition region between red and green dotted circles, the blended vector field evolves from source to vortex and then to sink as the distance from the singular point increases.
@@
The blended vector field defined in (10) satisfies the following properties:
@@
    The integral curves of $\mathbf{F}$ have continuous curvature bounded by $\bar{\kappa}$, if
$r_{i+1}-r_{i}\geq 3\rho,\quad i=1,2.$ (14a)
$r_{i}\geq r_{i+1}/2,\quad i=1,2.$ (14b)
@@
Remark 2
The physical meaning of the condition (14) is explained as follows. In $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$, the vector field gradually evolves from the source and sink to the vortex as it approaches the limit cycle. As the regions $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$ expand, the transition becomes smoother and the curvature in the blended region decreases. Condition (14a) allocates enough space for the blended vector field to steer its orientation with bounded curvature when approaching the limit cycle. Moreover, the domain of vortex should be expelled away from the singular point by increasing the region radii $r_{i}$, since the curvature of the vortex becomes unbounded around the singular point. The condition (14b) imposes a lower bound on the region radii such that the domain of vortex in the blended VF, i.e., $\mathcal{A}_{2}\bigcup\mathcal{A}_{3}$, is far enough from the singular point to avoid violation of the curvature constraint.
Since the stable limit cycle $\mathcal{O}_{2}$ is circular, we can always find a point $\bm{p}_{\theta_{d}}\in\mathcal{O}_{2}$ where the tangent vector aligns with the target orientation $\theta_{d}$, i.e., $\mathbf{F}(\bm{p}_{\theta_{d}})=[\cos\theta_{d},\sin\theta_{d}]^{\intercal}$. By translating the $\bm{p}_{\theta_{d}}$ to the target position $\bm{p}_{d}$, we establish the limit cycle as the target positive limit set and present the CVF as below:
@@
The curvature-constrained vector field given by (15) has the following properties:
@@
Figure 4: The proposed CVF $\mathbf{T}$, where the singular point is the origin.
The proposed CVF is illustrated by Figure 4. In the above theorem, the property (16a) means that the curvature of integral curves of the CVF is continuous and bounded by $\bar{\kappa}$ for any non-singular position. The property (16b) indicates that the integral curves of the CVF converge to a stable limit cycle containing the target position $\bm{p}_{d}$ almost globally. Moreover, the property (16c) ensures that the orientation of the CVF aligns with the target orientation $\theta_{d}$ at $\bm{p}_{d}$. By moving along the integral curves of the CVF, the nonholonomic robot approaches the target configuration while converging to the target positive limit set with continuous trajectory curvature bounded by $\bar{\kappa}$.
@@
Compared to existing VF-based motion planners with curvature constraints, the proposed CVF offers several key advantages. In [lau2015fluid], the VF only converges to the target position and depends on empirical parameter tuning. In contrast, our CVF guarantees that curvature-bounded integral curves simultaneously plan both position and orientation, as specified by the explicit condition in (14). While [qiao2024motion] also addresses position and orientation planning, it places the singularity of the curvature-constrained VF at the target position. Unlike this approach, our CVF establishes a stable limit cycle as the target positive limit set, keeping the target configuration away from the singular point and avoiding singularity during convergence. Although [pothen2017curvature] utilizes a stable limit cycle, the limited flexibility in its VF design makes it challenging to derive control inputs that both track the VF and maintain the curvature constraint. In contrast, our CVF offers more flexibility in parameter tuning, facilitating the co-design of the CVF and control laws in the subsequent section.
@@ IV Saturated Control Laws with Dynamic Gain
The CVF presented in Theorem 2 provides reference trajectories with desired curvature bound, converging to the target positive limit set. As pointed out in Section II-C, we will in this section find feasible control inputs of the nonholonomic robot, including the linear velocity $v_{x}$ and angular velocity $\omega$, to track the CVF under curvature constraint.
In accordance with the motion planning objective, we firstly specify the control law with user-defined bounds as follows:
@@
where $v_{\text{min}}$ is the lower bound, $c_{\bm{p}}$, $c_{\theta}$ are positive scalars, and the gain is given by $k_{v}=v_{\text{max}}-v_{\text{min}}$ such that the linear velocity is upper bounded by $v_{\text{max}}$. Moreover, $\theta_{e}$ is the orientation error between the robot and CVF defined by
@@
where the reference orientation $\theta_{r}$ of CVF is specified as
@@
Since the $\bm{p}_{\delta}$ is the singular point of the CVF, in the following we adopt the polar coordinates centered at $\bm{p}_{\delta}$:
$r_{\delta}$ $=\|\bm{p}-\bm{p}_{\delta}\|,$
$\varphi_{\delta}$ $=\arctan{\frac{\left<\bm{p}-\bm{p}_{\delta},\bm{e}_{y}\right>}{\left<\bm{p}-\bm{p}_{\delta},\bm{e}_{x}\right>}}.$
@@
It can be observed that the linear velocity $v_{x}$ decreases to the lower bound $v_{\text{min}}$ only if the robot reaches the target configuration $\bm{g}_{d}$, where $\bm{p}=\bm{p}_{d}$ and $\theta=\theta_{d}=\theta_{r}$. For robots such as wheeled UGV, $v_{\text{min}}$ could be chosen as zero so that the robot could asymptotically converge to $\bm{g}_{d}$. As for fixed wing UAVs and other robots with $v_{\text{min}}>0$, they are not able to converge to a specific configuration and thereby the result of convergence to the target positive limit set and repeated arrival at the target configuration is the best we can achieve.
When the robot’s orientation $\theta$ aligns with the reference orientation $\theta_{r}$, the robot’s trajectory becomes the time reparameterization of the integral curves of the CVF under the proposed linear velocity, referring to [chicone2006ordinary, Chapter 1.5]. This implies that the robot is able to track the CVF and converge to the target positive limit set by aligning its orientation with $\theta_{r}$ under the angular velocity:
@@
In the above equation, $k_{\omega}\theta_{e}$ is the feedback term and the feedforward term, i.e., the time derivative of $\theta_{r}$, is given by
@@
where $\Delta\theta=\theta-\theta_{\nabla}$, $\theta_{\nabla}$ and $A(r_{\delta})$ denote the direction and magnitude of the gradient $\nabla\theta_{r}$ of reference orientation, respectively. The detailed derivations of (21) are provided in Appendix D.
To ensure that the robot has curvature-bounded trajectory before it tracks the CVF, we refer to (2) and impose the angular velocity with a saturation bound given by
@@
and therefore, the saturated angular velocity is expressed as
@@
where $\omega_{0}$ is determined by(20), $\mathrm{sgn}(\cdot)$ is the sign function.
According to (45) and (46) in Appendix D, there holds that
@@
This equation implies that the upper bound of the feedforward term is larger than the saturation bound when $\bm{g}\in\mathcal{S}=\{\bm{g}\mid\|\bm{p}-\bm{p}_{\delta}\|<\rho\}$. Therefore, the angular velocity (22) possibly saturates in $\mathcal{S}$ even when the feedback term is omitted. In other words, $\mathcal{S}$ is the minimal region where saturation occurs with the presence of feedback term.
To stabilize the orientation error $\theta_{e}$, we aim to establish the condition under which the angular velocity (22) does not persistently saturate. In this way, $\theta_{e}$ is ultimately stabilized by $\omega=\omega_{0}$. From a geometric perspective, the robot’s trajectory, whose curvature is bounded by $\bar{\kappa}$, cannot be enclosed in an open circular area with radius $\rho=1/\bar{\kappa}$ such as $\mathcal{S}$. Enlightened by this, we ensure that (22) does not saturate persistently by confining the region where saturation occurs to the minimal extent of $\mathcal{S}$, as elaborated in the following theorem.
@@
The proposition that the angular velocity does not saturate persistently can be verified with a counter example. Assume that the angular velocity saturates all the time, which requires the robot to stay in $\mathcal{S}$. Since the target configuration $\bm{g}_{d}\notin\mathcal{S}$, the linear velocity satisfies $v_{x}>v_{min}\geq 0$. In this case, the robot moves along a circular arc with radius $\rho$ since $\left|\omega\right|=\bar{\omega}$ and $\kappa=\left|\bar{\omega}/v_{x}\right|=\bar{\kappa}=1/\rho$. However, such a circle cannot be contained within $\mathcal{S}$, as $\mathcal{S}$ is itself an open circular region with radius $\rho$. This implies that the robot will leave $\mathcal{S}$ even when its angular velocity persistently saturates, which contradicts with Theorem 3. Above discussions further imply that the robot leaves the region $\mathcal{S}$ and the saturation only occurs in finite time.
Due to the geometry of the CVF, i.e., the magnitude $A(r_{\delta})$ of gradient $\nabla\theta_{r}$ exceeds the maximum curvature around the singular point $\bm{p}_{\delta}$. Therefore, the saturation of the angular velocity within $\mathcal{S}$ is inevitable. To ensure that the robot leaves such region and hence the angular velocity only saturates in finite time, we propose the dynamic gain in the above theorem. However, this dynamic gain vanished when the robot circulates on the boundary of $\mathcal{S}$, i.e., $k_{\omega}=0$, $\forall\bm{g}\in\mathcal{N}_{k_{\omega}}=\left\{\bm{g}\mid r_{\delta}=\rho,\lvert\theta_{e}\rvert=\pi/2\right\}$. In addition, due to the singularity of the CVF and the definition of orientation error, there are two singular sets for the robot, i.e., $\mathcal{N}_{\mathbf{T}}=\left\{\bm{g}\mid\bm{p}=\bm{p}_{\delta}\right\}$ and $\mathcal{N}_{\theta}=\left\{\bm{g}\mid\lvert\theta_{e}\rvert=\pi\right\}$. The above singular sets together form the non-converging set $\mathcal{N}=\mathcal{N}_{\mathbf{T}}\cup\mathcal{N}_{\theta}\cup\mathcal{N}_{k_{\omega}}$ of the closed-loop system.
It is worth noting that the dynamic gain not only responds to the change in CVF under robot’s motion, it also imposes the inherent constraint $A(r_{\delta})\leq\bar{\kappa}$, $\forall r_{\delta}\geq\rho$ on the geometric properties of CVF. To address the above intrinsic coupling between the CVF and control laws under the curvature constraint and to ensure the existence of dynamic gain, the co-design of the CVF and control laws is of great necessity. Based on the flexible adjustment of the CVF’s parameters under (14), we establish the condition that ensures the orientation error is stabilized by the saturated control laws with dynamic gain in the following theorem.